Solve
a) log(x-2) + log(x+1) = 1
Can you please help me out? Thanks so much in advance:)

Answer by mananth(16949)

log(x-2) + log(x+1) = 1

=> log((x-2)(x+1))=log 10
=> (x-2)(x+1)=10
=> x^2-x-2=10
=> x^2-x-12=0
=>x^2-4x+3x-12=0
=>x(x-4)+3(x-4)=0
=> (x-4)(x+3)=0
x= 4 OR -3